4. A company estimates its manufacturing overhead will be $810,000 for the next year. What is...

4.

A company estimates its manufacturing overhead will be $810,000 for the next year. What is the predetermined overhead rate given the following independent allocation bases? When required, round your answers to nearest cent.

A. Budgeted direct labor hours: 54,000

$ ____per direct labor hour

B. Budgeted direct labor expense: $1,350,000

$___ per direct labor dollar

C. Estimated machine hours: 90,000

$___ per machine hour
Answer #1

A. Predetermined overhead rate = 810000/54000 = $15 per direct labor hour

B. Predetermined overhead rate = 810000/1350000 = $0.60 per direct labor dollar

C. Predetermined overhead rate = 810000/90000 = $9 per machine hour